The Scotia Company issues $100,000, 10% bonds at 102 on April 1, 2020. The bonds are dated January 1, 2020 and mature ten years from that date. Straight-line amortization is used. Interest is paid annually each December 31. Compute the bond carrying value as of December 31, 2027.

I keep receiving the wrong computation for this question! Can somebody please include the 3 months because this bond is issued on April 1st. Please please show the correct computation, thank you.
Answer #1

No of month for amortization = 10*12 = 120

Month from April 1,2020 to December 31,2027 = 7*12+9 = 93

Bond issue price = 100000*1.02 = 102000

Premium on bonds payable = 102000-100000 = 2000

Amortized premium = 2000*93/120 = 1550

Unamortized premium = 2000-1550 = 450

Carrying value on December 31,2027 = 100000+450 = $100450